Given the same reactant concentrations, the reaction

CO(g) + Cl2(g) →COCl2(g)

at 243°C is 97.5 times as fast as the same reaction at 153°C. Calculate the activation energy for this reaction in kj/mol.  Assume that the frequency factor is constant.
Answer #1

CO(g) + Cl2(g) →COCl2(g)

T1 = 153 + 273    = 426K

T2 = 243+273   = 516K

lnK2/K1    =     Ea/R[ 1/T1- 1/T2]

ln(97.5/1)     = Ea/8.314[1/426-1/516]

4.5798          = Ea(0.00234-0.00194)/8.314

Ea                 = 4.5798*8.314/0.0004   = 95191J      = 95.191KJ/mole